# Wildfire Risk Prediction: A Review

### v3

@@ 7.2.2 Transformers for Wildfire Risk Prediction
Transformers utilize multi-head self-attention mechanisms to calculate attention scores, enabling them to capture dependencies at any position within the sequence. The attention mechanism is defined as:
@@
Transformers operate on sequences transformed into tokens, which are then embedded with positional encoding to maintain the sequence order absent in traditional RNNs. The sequence of embedded tokens is processed by the encoder, which transforms it into a context vector using layers of multi-head self-attention and feed-forward networks. The decoder similarly processes the context vector to predict the output sequence.
The ability of Transformers to process sequence data without relying on sequential processing allows for parallel computation during training, significantly speeding up the learning process compared to RNNs (Vaswani et al., 2017). This makes them particularly effective for long-sequence predictions or complex scenarios requiring modeling of long-range dependencies (Zerveas et al., 2021; Prapas et al., 2023).
Transformers have been adapted for several predictive tasks in wildfire risk assessment.For example, Miao et al. (2023) developed a Transformer model with a window-based attention mechanism to predict forest fire risks in Chongli District, Beijing, using time-series data of meteorological, topographical, vegetation, and anthropogenic factors. The window mechanism confines attention to a fixed-size window centered on each element, reducing computational complexity and memory usage. Similarly, Cao et al. (2024) employed a Transformer to predict next-day wildfire risks in Quanzhou County, Guangxi Province, based on three days of meteorological, topographical, and human activity data. The model’s performance, analyzed using IGR, outperformed LSTM, RNN, and SVM in terms of generalization, noise resistance, and overfitting mitigation.
In addition, Prapas et al. (2023) introduced a TeleViT model that incorporates teleconnections to model long-range spatio-temporal interactions globally, aiming to predict subseasonal to seasonal wildfire patterns. This model combines fine-grained local-scale inputs with coarse-grained global-scale data, improving prediction accuracy for global wildfire patterns up to four months in advance.
@@
These examples illustrate the growing importance and versatility of Transformers in addressing the complexities of wildfire risk prediction, offering advancements over traditional models in handling long-range dependencies and large-scale data integration.
From the analysis above, it is evident that both RNNs and GRUs are less frequently used in wildfire risk prediction, which includes forecasting occurrences, susceptibility, and spread. In contrast, LSTM networks have a relatively higher usage rate. RNNs, compared to LSTMs, are particularly challenged by issues such as vanishing and exploding gradients and exhibit weaker capabilities in modeling long-distance dependencies. Furthermore, although some researchers argue that GRUs perform comparably to LSTMs (Viswanathan et al., 2019; Gao et al., 2020; Liu et al., 2021; Gao et al., 2021; Zarzycki and Ławryńczuk, 2022), the simplified structure of GRUs may result in inferior performance when processing highly complex sequences or capturing intricate relationships (Cahuantzi et al., 2023), especially in wildfire risk prediction tasks that do not require high real-time processing capabilities or are intended for edge computing devices. Recently, however, with the advent of Transformers, these models are gaining popularity in wildfire risk prediction due to their complex architecture, larger parameter capacity, and the advantages of parallel computation.
@@ 7.3 Image Semantic Segmentation and Classification Methods
Although RNNs and Transformers excel at capturing temporal dependencies for sequence-to-sequence learning tasks, these methods primarily classify individual pixels’ temporal sequence data. However, wildfire occurrences are not only influenced by local conditions but also by neighboring areas (e.g., wind propagation) (zhang2019forest). When RNNs use 1D data inputs, the spatial structure of the image is often lost (Hu et al., 2020). Consequently, several studies have explored the use of classical image semantic segmentation and classification methods, such as CNNs, GNNs, and Transformers, to predict wildfire risk at time $t$ using the influencing factors (e.g., fuel, weather) from time $t^{\prime}<t$. These methods rely on 2D image inputs, which explicitly model the spatial dependencies of input features, while also implicitly modeling the influence of current driving factors on future wildfire risk by having different input features and predicted time steps. In wildfire risk prediction, convolutional neural networks (CNNs) are commonly used, with graph neural networks (GNNs) and Transformers increasingly adopted in recent studies.
@@ 7.3.1 Convolutional Neural Networks
CNNs are foundational network structures in deep learning, designed to extract local spatial features through convolution operations, making them well-suited for capturing the spatial information necessary for wildfire risk prediction, as shown in Fig.10. Zhang et al. (2019) employed a modified AlexNet (Krizhevsky et al., 2012) using annual fire point maps and springtime mean influencing factors for Yunnan Province, China, to classify image patches as wildfire-prone areas. The model outputs a probability map of wildfire risk through softmax mapping in the final CNN layer. The authors compared the performance of the CNN model with RF, SVM, MLP, and kernel logistic regression models, finding that CNN achieved the highest overall accuracy. They also explored the contributions of various wildfire drivers using information gain ratio, identifying temperature, wind speed, surface roughness, precipitation, and elevation as the most influential factors.
@@
Subsequently, they extended the modified AlexNet to generate a global quarterly wildfire susceptibility map (Zhang et al., 2021), comparing its performance with MLP-2D, CNN-1D, and MLP-1D models. Results indicated that CNN-2D and MLP-1D models outperformed CNN-1D and MLP-2D. Additionally, the authors applied an improved permutation importance (PI) method and partial dependence plots (PDP) to investigate the interpretability of the CNN-2D model. PI analysis revealed that maximum monthly temperature, soil temperature, NDVI, and soil moisture were significant factors affecting the model. PDP analysis further indicated a negative relationship between cumulative precipitation and wildfire probability, while maximum monthly temperature and soil temperature positively influenced wildfire occurrence.
@@
Furthermore, they used a similar CNN method to predict the spatiotemporal variation of wildfire risk under different climate scenarios, discovering that global warming would lead to an increase in burned areas and a northward shift of wildfire-prone regions (Zhang et al., 2024a). Kanwal et al. (2023) similarly explored CNN-1D and CNN-2D classification models and machine learning algorithms for evaluating seasonal wildfire risks in Pakistan, finding that CNN-2D outperformed CNN-1D and other machine learning algorithms.
@@
Some studies have employed deeper or structurally and parameter-optimized CNN networks. For example, Oak et al. (2024) used VGG16 combined with ReLU, BatchNorm, Dropout, and a sigmoid layer to predict wildfire susceptibility in Quebec, Canada. Comparative experiments demonstrated that this model outperformed Xception, VGG16, ResNet50, and Inception V3. Similarly, Nur et al. (2022) applied a meta-heuristic optimization algorithm to fine-tune the hyperparameters of a CNN model for wildfire susceptibility mapping in the Plumas National Forest, USA. Their IGR analysis of 16 driving factors revealed that land use, drought index, and maximum temperature had the greatest impact on wildfire susceptibility. After applying the meta-heuristic optimization algorithm, CNN performance improved significantly. Likewise, (You et al., 2023) employed particle swarm optimization to optimize the structure and parameters of a CNN classification model to assess annual forest fire risk across China. Comparative analysis showed that this approach outperformed both CNN and traditional machine learning methods. Using SHAP for interpretability analysis, the most significant factors were found to be NDVI, land cover, and temperature. Eddin et al. (2023) introduced location information to the CNN by adding a separate convolution branch, highlighting the distinction between dynamic variables (those changing over time) and static variables (those that remain constant over the study period). To address the causal effects of static variables on dynamic ones, they introduced a LOAN layer to adjust the activation maps in the dynamic branch based on static features. Additionally, they used a sinusoidal encoding of Julian days to provide the model with explicit time information. Their experiments on the FireCube dataset, covering parts of the Mediterranean (Prapas et al., 2022; Kondylatos et al., 2023), demonstrated that the proposed model’s predictions of $t+1$ wildfire spread outperformed baseline models, including LSTM, ConvLSTM, RF, and XGBoost. Dong et al. (2024) proposed a similar architecture where dynamic and static variables are modeled separately in the encoding phase, resulting in improved accuracy in predicting fire radiative power.
Beyond classification-based approaches, some studies have adopted end-to-end models to provide higher-resolution predictions. For instance, Hodges and Lattimer (2019) proposed a deep convolutional inverse graphic network comprising convolution and deconvolution layers to simulate wildfire spread over homogeneous and heterogeneous landscapes for 6-hour and 24-hour intervals. The study demonstrated the high efficiency of CNNs in wildfire spread prediction. Similarly, Santopaolo et al. (2021) used an end-to-end CNN image segmentation model to predict wildfire risk in Sicily, Italy, and California, USA, based on driving factors from 8 days prior. Likewise, Bergado et al. (2021) designed an AllConvNet, a fully convolutional network, to predict wildfire burn probability for the next 7 days in Victoria, Australia, based on pre-fire driving factors. Overall, the proposed AllConvNet achieved higher prediction accuracy than logistic regression, SegNet, and MLP models. Feature extraction and normalization of LR coefficients revealed that total rainfall, lightning density, and surface temperature were among the most influential variables across all models. Additionally, Li et al. (2023b) introduced spatial and channel attention mechanisms to help the end-to-end CNN model focus on critical features while suppressing irrelevant ones to predict the next moment’s wildfire spread. Unlike previous studies, this model’s input also included the burning map from the prior time step. Comparative results demonstrated that this model outperformed other CNN models and ConvLSTM. Similarly, Shadrin et al. (2024b) employed the MA-Net segmentation model, which includes position attention and multi-scale attention blocks, to predict wildfire extent for the next 1-5 days. In contrast to previous studies, they also incorporated short-term weather forecasts for the next 1-5 days along with various wildfire driving factors. Experimental results showed accurate wildfire spread predictions within 3 days, with wind component, land cover, NDVI, EVI, and LAI being the most influential factors. Marjani et al. (2024a) combined CNN with atrous spatial pyramid pooling to model wildfires at different scales and shapes, adopting an end-to-end approach to model wildfire spread for the next day. They also used linear regression and Grad-CAM heatmap correlations to show that larger and smaller dilation rates help extract large-scale and small-scale wildfires, respectively. Similar to Eddin et al. (2023), Jiang et al. (2023) employed three branches to process static DEM and fuel data, dynamic wind speed, wind direction, temperature, and humidity data, as well as the previous time step’s wildfire status. The experiments demonstrated that this model’s inference efficiency was significantly higher than that of cellular automata and Farsite models. Furthermore, the model accurately predicted burned areas in real wildfire scenarios.
@@ 7.3.2 Other Methods
In addition to CNNs, other image segmentation methods, such as Graph Neural Networks (GNNs) and Transformers, have also been applied to wildfire risk prediction. GNNs, compared to CNNs, are better suited for handling non-Euclidean structured data (Asif et al., 2021), while Transformers’ multi-head self-attention mechanisms are more effective in modeling long-range semantic dependencies within images (Vaswani et al., 2017). For example, Jiang et al. (2022) proposed using an irregular graph network to simulate wildfire spread in variable-scale landscapes, addressing the issue of fixed-grid data being unable to adaptively differentiate landscape heterogeneity. The model adjusts graph node density based on terrain complexity to achieve uniformity in nodes and edges at varying scales. Comparative experiments with cellular automata and FARSITE demonstrated that the proposed model exhibits competitive accuracy and efficiency.
@@
Additionally, Chen et al. (2022) incorporated a GNN into the knowledge graph framework proposed by Ge et al. (2022) to enable automated learning and prediction of burn areas. Specifically, the model first constructs a knowledge graph in the form of triples, then uses RotaE to compute representation vectors for each entity. These entity vectors are associated with nodes in the graph, and the vectors are subsequently fed into the GNN to facilitate automatic updates. Finally, link prediction algorithms are employed to predict the burn areas associated with the recorded nodes. Experiments conducted in Portugal’s Montesinho Natural Park demonstrated that this algorithm significantly outperformed traditional machine learning algorithms such as RF, SVM, and MLP.
Moreover, Li and Rad (2024) introduced spatial attention mechanisms and focal modulation into a U-shaped encoder-decoder network based on Transformers, known as Swin Unet. This end-to-end image segmentation approach utilizes wildfire risk driving factors from the previous day to predict wildfire risk for the following day across the continental United States. Comparative experiments using the extended Next Day Wildfire Spread dataset (Huot et al., 2022) showed that this model outperforms baseline models, including the standard UNet and Swin Unet.
@@ 7.4 Spatiotemporal Prediction Methods
The previously discussed time series prediction methods and image segmentation/classification methods each have their strengths in capturing temporal and spatial dependencies, respectively. However, they are unable to simultaneously handle both spatial and temporal dependencies and capture the complex cross-spatiotemporal interaction patterns. For instance, Prapas et al. (2021) pointed out that focusing only on spatial scale modeling could overestimate a model’s real-world performance. To integrate spatial structure and time series data, many studies have attempted to combine CNN, GNN, and Transformer models with RNN, LSTM, and GRU methods. This transforms the task of time series prediction or localized optimization between consecutive wildfire states (Jiang et al., 2023) into spatiotemporal sequence prediction tasks. There are three primary approaches to achieving this combination: separated spatial and temporal modeling, coupled spatiotemporal modeling, and spatially enhanced coupled spatiotemporal modeling.

### v4

@@ 7.2.2 Transformers for Wildfire Risk Prediction
Transformers utilize multi-head self-attention mechanisms to calculate attention scores, enabling them to capture dependencies at any position within the sequence (Vaswani et al., 2017). The attention mechanism is defined as:
@@
Transformers operate on sequences transformed into tokens, which are then embedded with positional encoding to maintain the sequence order absent in traditional RNNs. The sequence of embedded tokens is processed by the encoder, which transforms it into a context vector using layers of multi-head self-attention and feed-forward networks. The decoder similarly processes the context vector to predict the output sequence. The ability of Transformers to process sequence data without relying on sequential processing allows for parallel computation during training, significantly speeding up the learning process compared to RNNs (Vaswani et al., 2017). This makes them particularly effective for long-sequence predictions or complex scenarios requiring modelling of long-range dependencies (Zerveas et al., 2021; Prapas et al., 2023).
Transformers have been adapted for several predictive tasks in wildfire risk prediction. For example, Miao et al. (2023) developed a Transformer model with a window-based attention mechanism to predict forest fire risks in Chongli District, Beijing, China, using time-series data of meteorological, topographical, vegetation, and anthropogenic factors. The window mechanism confines attention to a fixed-size window centered on each element, reducing computational complexity and memory usage. Similarly, Cao et al. (2024) employed a Transformer to predict next-day wildfire risks in Quanzhou County, Guangxi Province, China, based on three days of meteorological, topographical, and human activity data. The model’s performance, analyzed using IGR, outperformed LSTM, RNN, and SVM in terms of generalization, noise resistance, and overfitting mitigation.
@@
From the analysis above, it is evident that both RNNs and GRUs are less frequently used in wildfire risk prediction, which includes forecasting occurrences, susceptibility, and spread. In contrast, LSTM networks have a relatively higher usage rate. RNNs, compared to LSTMs, are particularly challenged by issues such as vanishing and exploding gradients and exhibit weaker capabilities in modeling long-distance dependencies. Furthermore, although some studies argue that GRUs perform comparably to LSTMs (Viswanathan et al., 2019; Gao et al., 2020; Liu et al., 2021a; Gao et al., 2021; Zarzycki and Ławryńczuk, 2022), the simplified structure of GRUs may result in inferior performance when processing highly complex sequences or capturing intricate relationships (Cahuantzi et al., 2023), especially in wildfire risk prediction tasks that do not require high real-time processing capabilities or are intended for edge computing devices. Recently, however, with the advent of Transformers, these models are gaining popularity in wildfire risk prediction due to their complex architecture, larger parameter capacity, and the advantages of parallel computation.
@@ 7.3 Image Semantic Segmentation and Classification Methods
Although RNNs and Transformers excel at capturing temporal dependencies for sequence-to-sequence learning tasks, these methods primarily classify individual pixels’ temporal sequence data. However, wildfire occurrences are not only influenced by local conditions but also by neighboring areas (e.g., wind propagation). When RNNs use 1D data inputs, the spatial structure of the image is often lost (Hu et al., 2020). Consequently, several studies have explored the use of image semantic segmentation and classification methods, such as CNNs, GNNs, and Transformers, to predict wildfire risk at time $t$ using the influencing factors (e.g., fuel, weather) from time $t^{\prime}<t$. These methods rely on 2D image inputs, which explicitly model the spatial dependencies of input features, while also implicitly modeling the influence of current driving factors on future wildfire risk by having different input features and predicted time steps. In wildfire risk prediction, CNNs are commonly used, with GNNs and Transformers increasingly adopted in recent studies.
@@ 7.3.1 Convolutional Neural Networks
CNNs are foundational network structures in deep learning, designed to extract local spatial features through convolution operations, making them well-suited for capturing the spatial information necessary for wildfire risk prediction, as shown in Fig.10. For example, Zhang et al. (2019) employed a modified AlexNet (Krizhevsky et al., 2012) using annual fire point maps and springtime mean influencing factors for Yunnan Province, China, to classify image patches as wildfire-prone areas. The model outputs a probability map of wildfire risk through softmax mapping in the final CNN layer. When comparing the CNN model’s performance with RF, SVM, MLP, and kernel logistic regression models, the authors found that CNN achieved the highest overall accuracy. They also analyzed the contributions of various wildfire drivers using the information gain ratio, identifying temperature, wind speed, surface roughness, precipitation, and elevation as the most influential factors.
@@
Building upon this, the authors extended the modified AlexNet to generate a global quarterly wildfire susceptibility map (Zhang et al., 2021). They compared its performance with MLP-2D, CNN-1D, and MLP-1D models, finding that CNN-2D and MLP-1D models outperformed CNN-1D and MLP-2D. Additionally, they applied an improved permutation importance (PI) method and partial dependence plots (PDP) to investigate the interpretability of the CNN-2D model. The PI analysis highlighted that maximum monthly temperature, soil temperature, NDVI, and soil moisture were significant factors affecting the model. Furthermore, PDP analysis revealed a negative relationship between cumulative precipitation and wildfire probability, while maximum monthly temperature and soil temperature positively influenced wildfire occurrence.
@@
Moreover, using a similar CNN approach, they predicted the spatiotemporal variation of wildfire risk under different climate scenarios, discovering that global warming would lead to an increase in burned areas and a northward shift of wildfire-prone regions (Zhang et al., 2024). In another study, Kanwal et al. (2023) explored CNN-1D and CNN-2D classification models along with other machine learning algorithms for evaluating seasonal wildfire risks in Pakistan, finding that CNN-2D outperformed CNN-1D and other machine learning algorithms.
@@
In addition, some studies have employed deeper or structurally optimized CNN networks. For instance, Oak et al. (2024) utilized VGG16 with ReLU, BatchNorm, Dropout, and a sigmoid layer to predict wildfire susceptibility in Quebec, Canada. Comparative experiments demonstrated that this model outperformed other architectures like Xception, VGG16, ResNet50, and Inception V3. Similarly, Nur et al. (2022) applied a meta-heuristic optimization algorithm to fine-tune the hyperparameters of a CNN model for wildfire susceptibility mapping in the Plumas National Forest, USA. Their IGR analysis revealed that land use, drought index, and maximum temperature had the greatest impact on wildfire susceptibility, and the model’s performance improved significantly after optimization. Likewise, (You et al., 2023) used particle swarm optimization to refine the structure and parameters of a CNN classification model for assessing annual forest fire risk across China, showing that this approach outperformed both standard CNN and traditional machine learning methods. Using SHAP for interpretability analysis, they identified NDVI, land cover, and temperature as the most significant factors.
Additionally, Eddin et al. (2023) introduced location information to the CNN by adding a separate convolution branch to differentiate between dynamic variables (those changing over time) and static variables (those that remain constant over the study period). To address the causal effects of static variables on dynamic ones, they incorporated a LOAN layer to adjust the activation maps in the dynamic branch based on static features. They also used sinusoidal encoding of Julian days to provide the model with explicit time information. Their experiments on the FireCube dataset, which covers parts of the Mediterranean (Prapas et al., 2022; Kondylatos et al., 2023), demonstrated that the proposed model’s predictions of $t+1$ wildfire spread outperformed baseline models, including LSTM, ConvLSTM, RF, and XGBoost. In a similar vein, Dong et al. (2024) proposed an architecture that models dynamic and static variables separately in the encoding phase, improving the accuracy in predicting fire radiative power.
Beyond classification-based methods, some studies have implemented end-to-end models for higher-resolution predictions. For example, Hodges and Lattimer (2019) proposed a deep convolutional inverse graphic network comprising convolution and deconvolution layers to simulate wildfire spread over homogeneous and heterogeneous landscapes for 6-hour and 24-hour intervals, demonstrating the high efficiency of CNNs in wildfire spread prediction. Similarly, Santopaolo et al. (2021) used an end-to-end CNN image segmentation model to predict wildfire risk in Sicily, Italy, and California, USA, based on driving factors from 8 days prior.
Furthermore, (Allaire et al., 2021) developed a hybrid architecture combining convolutional and fully connected layers to emulate the output of a numerical fire spread simulator based on the Rothermel model of fire propagation. Similarly, Bergado et al. (2021) designed an AllConvNet, a fully convolutional network, to predict wildfire burn probability for the next 7 days in Victoria, Australia. This model achieved higher prediction accuracy compared to logistic regression, SegNet, and MLP models. Feature extraction and normalization of logistic regression coefficients revealed that total rainfall, lightning density, and surface temperature were among the most influential variables.
Additionally, Li et al. (2023c) introduced spatial and channel attention mechanisms to enhance an end-to-end CNN model’s focus on critical features while suppressing irrelevant ones for predicting the next moment’s wildfire spread. This model also incorporated the burning map from the previous time step, outperforming other CNN models and ConvLSTM. Similarly, Shadrin et al. (2024b) used the MA-Net segmentation model, which includes position attention and multi-scale attention blocks, to predict wildfire extent for the next 1-5 days. Unlike previous studies, they included short-term weather forecasts along with various wildfire drivers. Experimental results showed that their model accurately predicted wildfire spread within 3 days, with the wind components, land cover, NDVI, EVI, and LAI being the most influential factors.
Lastly, Marjani et al. (2024a) combined CNN with atrous spatial pyramid pooling to model wildfires at different scales and shapes, adopting an end-to-end approach for next-day wildfire spread prediction. They used linear regression and Grad-CAM heatmap correlations to demonstrate how different dilation rates help extract both large-scale and small-scale wildfires. Similarly, Jiang et al. (2023) employed a three-branch network to process static DEM and fuel data, dynamic wind speed, temperature, humidity data, and the previous time step’s wildfire status. Their experiments showed that this model’s inference efficiency significantly surpassed that of cellular automata and Farsite models, accurately predicting burned areas in real wildfire scenarios.
@@ 7.3.2 Other Methods
In addition to CNNs, other image segmentation methods, such as GNNs and Transformers, have also been applied to wildfire risk prediction. GNNs, compared to CNNs, are better suited for handling non-Euclidean structured data (Asif et al., 2021), while Transformers’ multi-head self-attention mechanisms are more effective in modeling long-range semantic dependencies within images (Vaswani et al., 2017). For example, Jiang et al. (2022) proposed using an irregular graph network to simulate wildfire spread in variable-scale landscapes, addressing the issue of fixed-grid data being unable to adaptively differentiate landscape heterogeneity. The model adjusts graph node density based on terrain complexity to achieve uniformity in nodes and edges at varying scales. Comparative experiments with cellular automata and FARSITE demonstrated that the proposed model exhibits competitive accuracy and efficiency.
@@
Additionally, Chen et al. (2022) incorporated a GNN into the knowledge graph framework proposed by Ge et al. (2022) to enable automated learning and prediction of burn areas. Specifically, the model first constructs a knowledge graph in triples and uses RotaE to compute representation vectors for each entity. These entity vectors are associated with nodes in the graph, and the vectors are subsequently fed into the GNN to facilitate automatic updates. Finally, link prediction algorithms are employed to predict the burn areas associated with the recorded nodes. Experiments conducted in Portugal’s Montesinho Natural Park demonstrated that this algorithm significantly outperformed traditional machine learning algorithms such as RF, SVM, and MLP.
Moreover, Li and Rad (2024) introduced spatial attention mechanisms and focal modulation into a U-shaped encoder-decoder network based on Transformers, known as Swin Unet. This end-to-end image segmentation approach utilizes wildfire risk driving factors from the previous day to predict wildfire risk for the following day across the continental United States. Comparative experiments using the extended Next Day Wildfire Spread dataset (Huot et al., 2022) showed that this model outperforms baseline models, including the standard UNet and Swin Unet. In addition, Prapas et al. (2023) introduced a TeleViT model that incorporates teleconnections to model long-range spatial interactions globally, aiming to predict subseasonal to seasonal wildfire patterns. This model combines fine-grained local-scale inputs with coarse-grained global-scale data, improving prediction accuracy for global wildfire patterns up to four months in advance.
@@ 7.4 Spatiotemporal Prediction Methods
The previously discussed time series prediction methods and image segmentation/classification methods each have their strengths in capturing temporal and spatial dependencies, respectively. However, they are unable to simultaneously handle both spatial and temporal dependencies and capture the complex cross-spatiotemporal interaction patterns. For instance, Prapas et al. (2021) pointed out that focusing only on spatial scale modeling could overestimate a model’s real-world performance. To integrate spatial structure and time series data, many studies have attempted to combine CNN, GNN, and Transformer models with RNN, LSTM, and GRU methods. This transforms the task of time series prediction or spatially localized optimization between consecutive wildfire states (Jiang et al., 2023) into spatiotemporal sequence prediction tasks. There are three primary approaches to achieving this combination: separated spatial and temporal modeling, coupled spatiotemporal modeling, and spatially enhanced coupled spatiotemporal modeling.